OpenStudy (anonymous):

If the first step in the solution of the equation 2x - 8 = 5x + 3 is "subtract 2x," then what should the next step be?

pooja195 (pooja195):

\[\huge~\rm~2x−8=5x+3\]Subtract 5x from both sides. \[\huge~\rm~2x−8−5x=5x+3−5x\] \[\huge~\rm~−3x−8=3\] Add 8 to both sides. \[\huge~\rm~−3x−8+8=3+8\]

pooja195 (pooja195):

can you continue from that?

OpenStudy (anonymous):

yeah

pooja195 (pooja195):

let me know your answer

OpenStudy (anonymous):

kk